OpenStudy (anonymous):

Ben dives off a 7m platform. He reaches a maximum height of 7.3m after 0.27s. how long does it take him to reach the water ?

OpenStudy (cwrw238):

he falls under force of gravity use one of the equations of motion for constant acceleration g = 9.8 ms_2, s = height = 7.3 m, initial velocity = 0 so 7.3 = 0t + 0.5 * 9.8 * t^2 4.9t^2 = 7.3 t^2 = 4.9 / 7.3 find t and add 0.27 secs will give you total time in the air